# Synergistic Dynamical Decoupling and Circuit Design for Enhanced Algorithm Performance on Near-Term Quantum Devices

### v1

@@ III.1.3 Native gate sets
This section explores the performance of quantum algorithms implemented on two distinct sets of IBM QPUs. The first set comprises four 27-qubit devices utilizing the CX gate as their native two qubit gate, while the second set consists of four 127-qubit devices employing the ECR gate as their native two qubit gate.
@@
Figure 7: Comparison of two hardware native gate sets: {CX, ID, RZ, SX, X} and {ECR, ID, RZ, SX, X}, denoted as CX and ECR gate sets, respectively. Results obtained using four 27-qubit QPUs ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen for the CX gate set, and four 127-qubit QPUs ibm_kyoto, ibm_cusco, ibm_brisbane, and ibm_sherbrooke for the ECR gate set: (a) NAR, (b) NSP, (c) $\Delta_{\mathrm{NAR}}$, (d) $\Delta_{\mathrm{NSP}}$, (e) circuit fidelity, and (f) $\ln(\tau)$. The CPMG sequence is used for all data points. Each line represents a linear fit to the corresponding data.
@@
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{CX}}$ 0.548 $y=-0.056x+0.967$ $-0.832$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{CX}}$ 0.656 $y=-0.047x+1.010$ $-0.812$ 0
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{ECR}}$ 0.712 $y=-0.043x+1.033$ $-0.766$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{ECR}}$ 0.771 $y=-0.037x+1.052$ $-0.911$ 0
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{CX}}$ 0.464 $y=-0.068x+0.976$ $-0.678$ 0
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{CX}}$ 0.582 $y=-0.051x+0.967$ $-0.504$ 0.00092
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{ECR}}$ 0.618 $y=-0.056x+1.041$ $-0.579$ 0.00009
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{ECR}}$ 0.716 $y=-0.047x+1.072$ $-0.529$ 0.00044
@@
$\ln(\tau)$ (CX) 10.703 $y=\phantom{-}0.188x+9.294$ $8910$
$\ln(\tau)$ (ECR) 10.577 $y=\phantom{-}0.138x+9.543$ $7910$
Benchmark results obtained from the two QPU sets are presented in Fig. 7. Table 4 summarizes average values of metrics and parameters derived from the linear fits of these metrics against the number of qubits. As shown in Figs. 7(a) and 7(b), applying DD sequences generally improves NAR and NSP, respectively, for both native gate sets. The highest performance is achieved with $\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{ECR}}$ and $\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{ECR}}$, which leverages ECR-based QPUs and incorporates DD sequences. Moreover, QPUs utilizing ECR gate exhibit consistently higher baseline performance, $\mathrm{NAR}_{\mathrm{B}}^{\mathrm{ECR}}$ and $\mathrm{NSP}_{\mathrm{B}}^{\mathrm{ECR}}$, compared to those with CX gate, $\mathrm{NAR}_{\mathrm{B}}^{\mathrm{CX}}$ and $\mathrm{NSP}_{\mathrm{B}}^{\mathrm{CX}}$, even surpassing that with DD sequences, $\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{CX}}$ and $\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{CX}}$. This observation suggests that the selection of QPUs may be more critical for achieving optimal performance than relying solely on DD techniques. As shown in Figs. 7(c) and 7(d), DD effectiveness improves as the qubit count increases for both gate sets. Moreover, the CX gate set exhibits higher DD effectiveness than the ECR gate set. The circuit fidelity and $\ln(\tau)$ for each data point are presented in Figs. 7(e) and 7(f), respectively. ECR gate produces an overall higher circuit fidelity and lower schedule duration, potentially contributing to higher performance. The reported values of $\mathrm{EMSR}_{\mathrm{AR}}$ and $\mathrm{EMSR}_{\mathrm{SP}}$ are 92.5% and 72.5% for CX gate, respectively, resulting in an average 82.5% success rate. In comparison, these corresponding values are 75% and 62.5%, respectively, for the ECR gate, leading to an average 68.75% success rate. This suggests that DD sequences are more robust in mitigating errors for the CX gate set compared to the ECR gate set.
The correlation coefficient presented in Table 4 reveals a very strong native correlation between $\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{ECR}}$ and the number of qubits, compared to $\mathrm{NAR}_{\mathrm{B}}^{\mathrm{ECR}}$. Similarly, a very strong negative correlation is observed between circuit fidelity and qubit count for both CX and ECR gate sets, indicating a significant decrease in circuit fidelity as the number of qubits increases. Moreover, we observe a weaker correlation between NSP and qubit number compared to NAR, implying that the impact of qubit count on success probability is less pronounced than its effect on approximation ratio. Additionally, the high $p$-values for $\Delta_{\mathrm{NAR}}^{\mathrm{ECR}}$ and $\Delta_{\mathrm{NSP}}^{\mathrm{ECR}}$ suggest that the effectiveness of DD sequences for ECR gates is more susceptible to random fluctuations.
@@ III.2.1 Algorithm implementations
Figure 8: Comparison of CX and CZ implementations of QAOA across four 27-qubit IBM QPUs ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen. The CPMG sequence is used for all data points. Each line represents a linear fit of the data.
@@
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{CX}}$ 0.548 $y=-0.056x+0.967$ $-0.832$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{CX}}$ 0.656 $y=-0.047x+1.010$ $-0.812$ 0
$\mathrm{NAR}_{\mathrm{B}}^{\mathrm{CZ}}$ 0.638 $y=-0.049x+1.007$ $-0.850$ 0
$\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{CZ}}$ 0.691 $y=-0.047x+1.041$ $-0.885$ 0
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{CX}}$ 0.464 $y=-0.068x+0.976$ $-0.678$ 0
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{CX}}$ 0.582 $y=-0.051x+0.967$ $-0.504$ 0.00092
$\mathrm{NSP}_{\mathrm{B}}^{\mathrm{CZ}}$ 0.635 $y=-0.041x+0.940$ $-0.473$ 0.00208
$\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{CZ}}$ 0.653 $y=-0.056x+1.071$ $-0.589$ 0.00006
@@
$\Delta_{\mathrm{NSP}}^{\mathrm{CZ}}$ 0.018 $y=-0.015x+0.132$ $-0.202$ 0.21130
@@
$\ln(\tau)$ (CX) 10.703 $y=\phantom{-}0.188x+9.294$ $8910$
$\ln(\tau)$ (CZ) 10.799 $y=\phantom{-}0.192x+9.360$ $9010$
We compare the performance of CX and CZ implementations of QAOA on four 27-qubit IBM QPUs, as detailed in Sec. II.3. The CPMG sequence is consistently utilized throughout this analysis. The results are presented in Fig. 8. As shown in Figs. 8(a) and 8(b), CZ implementation with DD sequences achieves the highest average values for both NAR ($\mathrm{NAR}_{\mathrm{DD}}^{\mathrm{CZ}}$) and NSP ($\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{CZ}}$). However, for NSP at qubit counts exceeding 9, CZ implementation without DD sequences, $\mathrm{NSP}_{\mathrm{B}}^{\mathrm{CZ}}$, outperforms that with DD sequences, $\mathrm{NSP}_{\mathrm{DD}}^{\mathrm{CZ}}$, potentially due to the introduction of significant errors by DD sequences themselves. Furthermore, CX implementation exhibits an increasing DD effectiveness $\Delta_{\mathrm{NAR}}$ and $\Delta_{\mathrm{NSP}}$, as the qubit number grows (Figs. 8(c) and 8(d)). Conversely, $\Delta_{\mathrm{NSP}}$ for CZ implementation, $\Delta_{\mathrm{NSP}}^{\mathrm{CZ}}$, slightly decreases with increasing qubit count. Additionally, $\mathrm{EMSR}_{\mathrm{AR}}$ and $\mathrm{EMSR}_{\mathrm{SP}}$ are consistently higher for CX implementation (92.5% and 80%, respectively) compared to CZ implementation (75% and 57.5%, respectively). This translates to an average EMSR of 86.25% for CX implementation and 66.25% for CZ implementation, suggesting higher robustness of DD sequences for CX implementation. Circuit fidelity and schedule duration are shown in Figs. 8(e) and 8(f), respectively. Both CX and CZ implementations exhibit comparable circuit fidelities. For a larger number of qubits, CX implementation even achieves slightly higher fidelities. Moreover, CX implementation demonstrates a consistently shorter schedule duration compared to CZ implementation. This difference in schedule duration is attributed to the increased number of single-qubit gates required by CZ implementation (details in Fig. 3).
Table 5 summarizes the average value of each metric along with the linear fit parameters extracted from Fig. 8. The negative coefficient associated with $\Delta_{\mathrm{NSP}}^{\mathrm{CZ}}$ suggests a decrease in DD effectiveness with an increasing qubit count for CZ implementation. Furthermore, correlation coefficients between NAR ($\mathrm{NAR}_{\mathrm{B}}$ and $\mathrm{NAR}_{\mathrm{DD}}$) and qubit count reveal that applying DD sequences weakens the correlation for CX implementation while strengthening it for CZ implementation. This trend is also observed for $\mathrm{NSP}_{\mathrm{B}}$ and $\mathrm{NSP}_{\mathrm{DD}}$. An additional observation is the high $p$-values associated with $\Delta_{\mathrm{NAR}}^{\mathrm{CZ}}$ and $\Delta_{\mathrm{NSP}}^{\mathrm{CZ}}$ indicating the potential dominance of random fluctuations in these metrics for CZ implementation.
The results suggest that although CX implementation offers more advantages in terms of DD effectiveness, the higher performance of CZ implementation highlights the significance of circuit structure in executing QAOA. In certain instances, the inherent advantage of a more symmetrical circuit structure, as exhibited by CZ implementation, can outweigh the benefits of strong DD mitigation achieved with CX implementation. However, the optimal selection of gate decomposition ultimately relies on the specific algorithm being implemented, the hardware capabilities available, and the desired balance between overhead caused by DD and potential performance gains.

### v2

@@ III.1.3 Native Gate Sets
This section explores the performance of quantum algorithms implemented on two distinct sets of IBM QPUs. The first set comprises four 27-qubit devices utilizing the CX gate as their native two-qubit gate, while the second set consists of four 127-qubit devices employing the ECR gate as their native two-qubit gate.
@@
Figure 7: Comparison of two hardware-native gate sets: {CX, ID, RZ, SX, X} and {ECR, ID, RZ, SX, X}, denoted as CX and ECR gate sets, respectively. Results obtained using four 27-qubit quantum processing units (QPUs) ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen for the CX gate set, and four 127-qubit QPUs ibm_kyoto, ibm_cusco, ibm_brisbane, and ibm_sherbrooke for the ECR gate set: (a) NAR, (b) NSP, (c) $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, (d) $\Delta_{\mathrm{\scriptscriptstyle NSP}}$, (e) circuit fidelity, and (f) $\ln(\tau/\mathrm{dt})$. The CPMG sequence is used for all data points. Each line represents a linear fit to the corresponding data.
Benchmark results obtained from the two QPU sets are presented in Fig. 7. Table 4 summarizes average values of metrics and parameters derived from the linear fits of these metrics against the number of qubits. As shown in Figs. 7(a) and (b), applying DD sequences generally improves the NAR and NSP, respectively, for both native gate sets. The highest performance is achieved with $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle ECR}}$ and $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle ECR}}$, which leverages ECR-based QPUs and incorporates DD sequences. Moreover, QPUs utilizing the ECR gate exhibit consistently higher baseline performance, $\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle ECR}}$ and $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle ECR}}$, compared to those with the CX gate, $\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CX}}$ and $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CX}}$, even surpassing those with DD sequences, $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CX}}$, and $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CX}}$. This observation suggests that the selection of QPUs may be more critical for achieving optimal performance than relying solely on DD techniques. As shown in Figs. 7(c) and (d), DD effectiveness improves as the qubit count increases for both gate sets. Moreover, the CX gate set exhibits higher DD effectiveness than the ECR gate set. As illustrated in Figs. 7(e) and (f), the ECR gate produces an overall higher circuit fidelity and lower schedule duration, potentially contributing to higher performance. Analyzing the proportion of positive outcomes in the experimental data presented in Figs. 7(c) and (d), we observe that the reported values of $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ and $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$ are 92.5% and 72.5% for the CX gate, respectively, whereas the corresponding values are 75% and 62.5%, respectively, for the ECR gate, suggesting that DD sequences are more robust in mitigating errors for the CX gate set.
The correlation coefficient presented in Table 4 reveals a very strong negative correlation between $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle ECR}}$ and the number of qubits compared to $\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle ECR}}$. Similarly, a very strong negative correlation is observed between circuit fidelity and qubit count for both CX and ECR gate sets, indicating a significant decrease in circuit fidelity as the number of qubits increases. Moreover, we observe a weaker correlation between NSP and qubit number compared to NAR, implying that the impact of qubit count on success probability is less pronounced than its effect on approximation ratio. Additionally, the high $p$-values for $\Delta_{\mathrm{\scriptscriptstyle NAR}}^{\mathrm{\scriptscriptstyle ECR}}$ and $\Delta_{\mathrm{\scriptscriptstyle NSP}}^{\mathrm{\scriptscriptstyle ECR}}$ suggest that the effectiveness of DD sequences for ECR gates is more susceptible to random fluctuations. This can be attributed to the interplay between the intended decoherence suppression capabilities of DD sequences and the additional gate errors that they introduce. As shown in Figs. 7(e) and (f), circuits utilizing ECR gates demonstrate higher fidelities while requiring shorter execution times. The inherent advantage of shorter circuits may limit the potential for further enhancement through the use of DD sequences. In such cases, incorporating extra DD pulses could even lead to a decrease in overall algorithm performance.
@@
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CX}}$ 0.548 $y=-0.056x+0.967$ $-0.832$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CX}}$ 0.656 $y=-0.047x+1.010$ $-0.812$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle ECR}}$ 0.712 $y=-0.043x+1.033$ $-0.766$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle ECR}}$ 0.771 $y=-0.037x+1.052$ $-0.911$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CX}}$ 0.464 $y=-0.068x+0.976$ $-0.678$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CX}}$ 0.582 $y=-0.051x+0.967$ $-0.504$ 0.00092
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle ECR}}$ 0.618 $y=-0.056x+1.041$ $-0.579$ 0.00009
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle ECR}}$ 0.716 $y=-0.047x+1.072$ $-0.529$ 0.00044
@@
$\ln(\tau/\mathrm{dt})$ (CX) 10.703 $y=\phantom{-}0.188x+9.294$ $8910$
$\ln(\tau/\mathrm{dt})$ (ECR) 10.577 $y=\phantom{-}0.138x+9.543$ $7910$
@@ III.2.1 Algorithm Implementations
Figure 8: Comparison of CX and CZ implementations of QAOA across four 27-qubit IBM QPUs ibmq_mumbai, ibmq_kolkata, ibm_cairo, and ibmq_ehningen: (a) NAR, (b) NSP, (c) $\Delta_{\mathrm{\scriptscriptstyle NAR}}$, (d) $\Delta_{\mathrm{\scriptscriptstyle NSP}}$, (e) circuit fidelity, and (f) $\ln(\tau/\mathrm{dt})$. The CPMG sequence is used for all data points. Each line represents a linear fit of the data.
We compare the performance of CX and CZ implementations of the QAOA on four 27-qubit IBM QPUs, as detailed in Section II.3. The CPMG sequence is consistently utilized throughout this analysis. As shown in Figs. 8(a) and (b), CZ implementation with DD sequences achieves the highest average values for both the NAR ($\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CZ}}$) and NSP ($\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CZ}}$). However, for the NSP at qubit counts exceeding 9, CZ implementation without DD sequences, $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CZ}}$, outperforms that with DD sequences, $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CZ}}$, potentially due to the introduction of significant errors by DD sequences themselves. Furthermore, CX implementation exhibits increasing DD effectiveness as the qubit number grows (Figs. 8(c) and (d)), whereas $\Delta_{\mathrm{\scriptscriptstyle NSP}}$ for CZ implementation exhibits a slight decrease. Additionally, $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle AR}}$ and $\mathrm{EMSR}_{\mathrm{\scriptscriptstyle SP}}$ are consistently higher for CX implementation (92.5% and 80%, respectively) compared to CZ implementation (75% and 57.5%, respectively), suggesting a higher robustness of DD sequences for CX implementation. As depicted in Figs. 8(e) and (b), CX and CZ implementations exhibit comparable circuit fidelities. For a larger number of qubits, CX implementation even achieves slightly higher fidelities. Moreover, CX implementation demonstrates a consistently shorter schedule duration compared to CZ implementation (Fig. 8(f)). This difference in schedule duration is attributed to the increased number of single-qubit gates required by CZ implementation (details in Fig. 3).
Table 5 summarizes the average value of each metric along with the linear fit parameters extracted from Fig. 8. The negative coefficient associated with $\Delta_{\mathrm{\scriptscriptstyle NSP}}^{\mathrm{\scriptscriptstyle CZ}}$ suggests a decrease in DD effectiveness, as measured by success probability, with an increasing qubit count for CZ implementation. This behavior can be attributed to the intricate interplay between the optimization landscape and gate errors. As the number of qubits involved increases, the cumulative error introduced by DD sequences becomes more pronounced. This significantly alters the optimization landscape, rendering the pre-optimized parameters of the QAOA in the noiseless case no longer suitable. Furthermore, correlation coefficients between the NAR ($\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}$ and $\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}$) and qubit count reveal that applying DD sequences weakens the correlation for CX implementation while strengthening it for CZ implementation. This trend is also observed for $\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}$ and $\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}$. An additional observation is the high $p$-values associated with $\Delta_{\mathrm{\scriptscriptstyle NAR}}^{\mathrm{\scriptscriptstyle CZ}}$ and $\Delta_{\mathrm{\scriptscriptstyle NSP}}^{\mathrm{\scriptscriptstyle CZ}}$, indicating the potential dominance of random fluctuations in these metrics for CZ implementation.
The results suggest that although CX implementation offers more advantages in terms of DD effectiveness, the higher performance of CZ implementation highlights the significance of circuit structure in executing the QAOA. In certain instances, the inherent advantage of a more symmetrical circuit structure, as exhibited by CZ implementation, can outweigh the benefits of strong DD mitigation achieved with CX implementation. However, the optimal selection of gate decomposition ultimately relies on the specific algorithm being implemented, the hardware capabilities available, and the desired balance between overhead caused by DD and potential performance gains.
@@
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CX}}$ 0.548 $y=-0.056x+0.967$ $-0.832$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CX}}$ 0.656 $y=-0.047x+1.010$ $-0.812$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CZ}}$ 0.638 $y=-0.049x+1.007$ $-0.850$ 0
$\mathrm{NAR}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CZ}}$ 0.691 $y=-0.047x+1.041$ $-0.885$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CX}}$ 0.464 $y=-0.068x+0.976$ $-0.678$ 0
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CX}}$ 0.582 $y=-0.051x+0.967$ $-0.504$ 0.00092
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle B}}^{\mathrm{\scriptscriptstyle CZ}}$ 0.635 $y=-0.041x+0.940$ $-0.473$ 0.00208
$\mathrm{NSP}_{\mathrm{\scriptscriptstyle DD}}^{\mathrm{\scriptscriptstyle CZ}}$ 0.653 $y=-0.056x+1.071$ $-0.589$ 0.00006
@@
$\Delta_{\mathrm{\scriptscriptstyle NSP}}^{\mathrm{\scriptscriptstyle CZ}}$ 0.018 $y=-0.015x+0.132$ $-0.202$ 0.21130
@@
$\ln(\tau/\mathrm{dt})$ (CX) 10.703 $y=\phantom{-}0.188x+9.294$ $8910$
$\ln(\tau/\mathrm{dt})$ (CZ) 10.799 $y=\phantom{-}0.192x+9.360$ $9010$